OpenStudy (anonymous):

Can someone help me with this. I am using the formula: R=Squareroot(kQ/E) where Q is 1.5x10^-9 and E is 5.8x10^5 and K is given as 9x10^9 I am not getting the correct answer at all. How do I get it? A ball has an electric charge of +1.5 × 10^-9 coulombs. At what distance from the ball's center is the electric field strength equal to 5.8 × 10^5 newtons/coulomb? (k = 9.0 × 10^9 newton·meter^2/coulomb^2)

OpenStudy (abhisar):

Hello @cmorse ! What answer are you getting ?

OpenStudy (abhisar):



OpenStudy (abhisar):

\(\sf \large E\times4\pi r^2=\frac{Q}{\epsilon_o }\\->\huge \cancel{4\pi} r^2=\frac{\cancel{4\pi} .Q.k}{E}\)

OpenStudy (abhisar):

Plug in the values and find the value of r

OpenStudy (anonymous):

Then answer I have is 1.5x10^-2

OpenStudy (abhisar):

\(\huge \sf \frac{1.5\times10^{-9}\times9\times10^9}{5.8\times10^5}=\)

OpenStudy (abhisar):

@ash2326 can u check it ?

OpenStudy (anonymous):

Yep I got have done it that way and I still get an answer that is different than the one in the answer key.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar @kropot72 can you see it ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Can you see what?

OpenStudy (abhisar):

I am not sure about ur answer...that's why asking others.

OpenStudy (anonymous):

Oh okay. To be honest, I am not sure about it either. I am checking the answer key to make sure it is correct and I think this one is wrong also.

OpenStudy (abhisar):

i think the answer in answer key is wrong

OpenStudy (anonymous):

Okay, thank you. That is what I needed.

OpenStudy (anonymous):

4.82 mm

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @RaphaelFilgueiras 4.82 mm \(\color{blue}{\text{End of Quote}}\) yes that's what i am also getting ! So the answer key is wrong @cmorse